OpenStudy (anonymous):

a gas has a pressure of 120 kpa, a temperature of 400 k and a volume of 50 milliliters. what volume will the gas have at a pressure of 60 kpa and a temperature of 200 kpa?

OpenStudy (jfraser):

The ideal gas equation says\[P*V = n*R*T\] since you're not given moles, they must remain constant between the 2 states, so \[R = \frac{P*V}{T}\]Since R must remain constant under all conditions, then any other combination of P,V, and T must also equal the same R\[\frac {P_{1}*V_{1}}{T{_1}} = \frac{P_{2}*V{_2}}{T{_2}}\] You know all but 1 value, solve for V2.